# Repository-level Code Translation Benchmark Targeting Rust

### v3

@@ Abstract.
Recent advances in large language models (LLMs) have shown significant capabilities in code translation, often evaluated using benchmarks like CodeTransOcean. However, these evaluations typically focus on simple, function-level translations without considering dependencies, which does not reflect the complexities of real-world software development. Further, their effectiveness in translating to newer, lower-resource languages like Rust in realistic scenarios is still under-explored.
To address this gap, we introduce first repository-level code translation benchmark comprising 375 tasks targeting Rust, complete with relevant dependencies. Using this benchmark, we study four state-of-the-art LLMs, analyzing their erroneous outputs to understand their performance in more complex translation scenarios. Our findings reveal that LLMs exhibit substantially worse performance (41.5%-56.2% Pass@1 drop of GPT-4) on repository-level translations compared to simpler tasks, highlighting limitations in existing evaluation methods. The model that performed the best is Claude-3.5, demonstrating the strongest translation capabilities in both basic functionality accuracy and several relevant additional abilities. Additionally, we discover that LLMs struggle with identifying language differences in complex tasks, and that increased dependencies correlate with greater translation difficulty.
11footnotetext: Mingwei Liu is the corresponding author.
Repository-level Code Translation, Rust, Large Language Models,
ccs: Software and its engineering Software notations and tools
@@ 1. Introduction
Code translation, also known as code migration, refers to the process of porting a project from one programming language to another. This process is often driven by the need to adapt to different runtime environments, optimize performance, or enhance security (nee, 2018, 2020b, 2020c, 2020d, 2020a; Thönes, 2015). In today’s rapidly evolving technological landscape, with the continuous emergence of new programming languages and architectural paradigms, such as Rust and Cangjie, there is a substantial demand for effective and efficient code translation. Efficiently migrating legacy code to new languages and environments remains a significant challenge for developers and organizations alike (Mateus et al., 2023; Kontogiannis et al., 2010).
In recent years, large language models have been increasingly utilized for tasks such as code generation and code translation, demonstrating promising results in translating code between widely used languages like Java and Python (Pan et al., 2024, 2023). However, their effectiveness in translating to newer, lower-resource languages like Rust is still under-explored. Rust presents unique challenges due to its limited training data and significant language-specific features that differ from those found in more established programming languages(rus, 2024). Consequently, the translation performance and generalization capabilities of large models for Rust might not be optimal.
Existing code translation datasets do include some coverage of Rust, but they tend to be overly simplistic and fail to capture the real-world demands of code translation in actual development scenarios (Yan et al., 2023). Most existing datasets comprise pairs of source and target language code (usually at the function level(Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023)) and accompanying unit tests to verify functional equivalence post-translation. However, the tasks used in current benchmarks often lack complexity, predominantly involving standalone functions without external dependencies (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a). Even the datasets that do extend to the file level generally only add minimal context to make the functions executable in isolation. Moreover, the sources of these datasets are typically online coding platforms or artificially constructed examples, which significantly diverge from real-world software development practices (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a; Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023; Yan et al., 2023). In real-world scenarios, developers often maintain the original project architecture and incrementally translate individual functions while relying on the repository’s context, which includes complex, inconsistent, and cross-file dependencies. These are aspects current benchmarks fail to address but are critical in real-world development.
To fill these gap, we propose the RustRepoTrans, a repository-level code translation benchmark targeting Rust. This benchmark was constructed by curating 375 code translation tasks that include relevant dependencies, reflecting the complexities found in actual software development projects.
@@
In our study, we investigate the performance of state-of-the-art LLMs on RustRepoTrans, focusing on their effectiveness in translating code with complex dependencies. We also assess how our benchmark presents greater challenges compared to existing benchmarks, analyzing the factors that influence task difficulty and the types of errors encountered by the models. Additionally, we explore essential capabilities that LLMs demonstrate beyond mere translation accuracy, such as noise robustness, syntactical difference identification and code simplicity.
Through our experiments, we find that existing LLMs struggle significantly with repository-level code translation compared to standalone tasks, the proportion of compilation errors reached as high as 94.8%, indicating that current evaluations do not adequately reflect their capabilities in real-world scenarios. We identify key challenges related to dependency handling and highlight that the complexity of generated code is indicative of translation effectiveness. Errors caused by function and variable dependencies accounted for a total of 61.9%. Overall, our findings provide critical insights into the limitations of LLMs in complex code translation tasks, paving the way for future improvements in model design and evaluation methodologies.
@@
The contributions of this paper are summarized as follows.
@@
  • Introduction of RustRepoTrans: We present the first repository-level code translation benchmark, consisting of 375 tasks targeting Rust, which includes relevant dependencies, providing a more realistic evaluation framework.
@@
  • Evaluation of LLMs in Complex Scenarios: We conduct a comprehensive study of four state-of-the-art LLMs on this benchmark, revealing their performance in more challenging code translation tasks that reflect real-world software development scenarios.
  • Error Analysis and Insights: We analyze the errors generated by the LLMs and categorize them into 10 error causes, offering insights into their limitations and the factors affecting their translation capabilities, particularly in handling dependencies.
@@ 2.1. Code Translation
Code translation is a crucial tool in modern software development, enabling interoperability among different programming languages and facilitating code reuse. It is particularly important for updating legacy systems and integrating components written in incompatible languages. Studies show that code translation improves productivity by reducing the time and cost of manual translation (Lu et al., 2021; Puri et al., 2021), while also aiding comprehension and maintenance in multilingual projects.
Historically, code translation relied on rule-based systems that used handcrafted syntactic and semantic rules. Early statistical machine translation (SMT) models improved accuracy by mapping frequent code patterns, as seen in datasets like Java-C and CoffeeScript-JavaScript (Nguyen et al., 2013; Chen et al., 2018). As research progressed, small neural networks and sequence models began to replace rule-based approaches, learning patterns from smaller datasets. However, early neural models, such as recurrent neural networks (RNNs), struggled with complex syntax and language-specific idioms (Dong and Lapata, 2016; Yin and Neubig, 2017, 2018).
Despite advances with large language models (LLMs), challenges persist, including incorrect translations and limited adaptability to coding styles (Zhu et al., 2022a; Rithy et al., 2022). This has led to increased focus on evaluating translation quality and the impact of prompt engineering. Jiao et al. (Jiao et al., 2023) developed the G-TransEval benchmark to assess translation complexity, revealing that models perform well on simple tasks but struggle with more complex ones. Similarly, Pan et al. (Pan et al., 2024) found that context-rich prompts can enhance translation reliability by an average of 5.5%. Prompt engineering strategies, such as those proposed by Yang et al. (Yang et al., 2024), utilize test cases to improve accuracy, particularly in challenging language pairs. Macedo et al. (Macedo et al., 2024a) emphasized the importance of consistent output formatting for accurate evaluation, advocating for controlled prompts and post-processing to ensure reliable benchmarking.
@@ 2.2. Code Translation Benchmarks
To address these gaps, some datasets, such as CoST (Zhu et al., 2022b) and CodeNet (Puri et al., 2021), have aimed to enhance language diversity and quality. However, issues like erroneous solutions and limited support for Rust persist. The latest benchmark, CodeTransOcean (Yan et al., 2023), expands coverage to 45 languages but still lacks adequate support for Rust, hindering comprehensive research in this area.
@@
RustRepoTrans is the first repository-level code translation benchmark focusing on Rust, designed to account for complex dependencies and derived from real-world projects, distinguishing it from previous benchmarks.
@@ 3. RustRepoTrans Benchmark
We begin by introducing the benchmark format, followed by the construction method of RustRepoTrans, and finally, we present the resulting benchmark.
@@ 3.1. Benchmark Format
Each code translation task in RustRepoTrans consists of a pair of functions along with their relevant dependencies, formatted as <source function, target function, target function dependencies, target function test cases>, as shown in Fig. 1. The function pairs <source function, target function> represent functionally equivalent code snippets from the source and target languages, along with their respective file paths. The target function dependencies include elements such as function dependencies, data type dependencies, variable dependencies, and library dependencies related to the target function.
In this benchmark, the source function, target function signature, and associated dependencies serve as input, with the goal for LLMs to generate a complete target function. The correctness of this generated function can then be verified using the associated test cases.
@@ 3.2. Benchmark Construction Method
Migration Project Selection. In this step, we select projects that have been rewritten in Rust from other languages, specifically C++, Java, and Python, due to their popularity and likelihood of having such rewritten versions. These projects are expected to contain functionally equivalent function pairs between their Rust and original language versions. We identify suitable projects by searching GitHub for terms like “Rust version” and “implemented in Rust”. focusing on larger projects (applying qualifier “size:>1000”) to ensure sufficient function pairs. After locating these projects, we verify their versions through documentation and code review. In this way, we identify pairs of source project and target project.
@@
Figure 2. The Prompt Used for Equivalent Function Pair Identification
@@ 3.3. Resulting Benchmark
This process enables us to construct RustRepoTrans, which comprises 375 tasks for repository-level code translation. Detailed information is provided in Table 1, alongside comparisons with other existing benchmarks. The size of RustRepoTrans is comparable to most previous benchmarks, and each task includes a manually verified ground truth for translation and corresponding unit tests, achieving over 90% average test coverage, indicating a high quality.
@@
Table 1. Comparison of Different Datasets for Code Translation.
@@
RustRepoTrans Github Function Level 150.1 375 C++, Java, Python
@@
Repository-level Dependency. RustRepoTrans focuses on code translation tasks with rich repository context, setting it apart from previous benchmarks. Unlike artificially constructed data or data sourced from Q&A websites, real project data exhibits more complex dependency relationships, including function, data type, and variable dependencies. Table 1 presents statistics on the presence of dependencies in existing datasets, revealing that only RustRepoTrans includes these crucial elements. This inclusion makes RustRepoTrans a more realistic benchmark, suitable for evaluating LLMs that must account for intricate file-level interactions and contextual dependencies.
@@
We analyzed 54 Rust keywords from the official documentation, comparing occurrences in CodeTransOcean(shown in Fig. 5) and RustRepoTrans (shown in Fig. 5) datasets. CodeTransOcean has 12 uncovered keywords, while RustRepoTrans has only 8, including critical asynchronous programming keywords async and await, which appear in 25.1% of the data. In term of keyword proportion differences between CodeTransOcean and RustRepoTrans (shown in Fig. 5), twelve keywords show negative differences, but only three exceed 6%: extern, let, and use. This is due to RustRepoTrans’s function-level data, resulting in 0% for library-related keywords like extern and use. Additionally, RustRepoTrans ’s real-world focus leads to less frequent use of the variable declaration keyword let compared to CodeTransOcean.
Figure 3. the proportion of data in the CodeTransOcean that contains each keyword.
Figure 4. the proportion of data in the RustRepoTrans that contains each keyword
Figure 5. the difference in the proportion of data containing keywords between RustRepoTrans and CodeTransOcean.
@@ 4.1. Experimental Setup
Model Selection. We selected four state-of-the-art LLMs that have been widely studied in recent code translation research or were recently released and have demonstrated strong performance on benchmark leaderboards (e.g.,SuperCLUE [^1] https://www.superclueai.com/ (Xu et al., 2023)). Table 2 provides an overview of the four LLMs used in our experiments, along with their release dates (Column “Time”) and model sizes (Column “Size”). As shown in Table 2, our study includes both open-source and closed-source LLMs, as well as models designed for general-purpose and code-specific applications. Due to resource constraints, we selected open-source models with sizes below 20B parameters. Note that our primary contribution is a new benchmark designed to provide a more challenging and realistic evaluation of LLMs’ code translation capabilities, closely simulating real-world development scenarios. Expanding the scope to include a larger and more diverse set of models is reserved for future work.
Implementation Details. For open-source models, we obtained and executed the released versions from their official repositories based on provided documentation. We used greedy sampling (Chen et al., 2015) as our generation strategy, generating a single solution code sample per task using greedy decoding (i.e.,setting the “do_sample” hyperparameter to false). All evaluations were conducted on an NVIDIA A800 80GB GPU. For closed-source LLMs, we accessed each model through its respective API interface (OpenAI, 2024; Claude, 2024) as of September 2024. To achieve results similar to greedy decoding, we set the “temperature” hyperparameter to 00. Since some of our data contains extensive dependencies, resulting in very long prompts, we set the “max_tokens” hyperparameter to the upper limit supported by all studied LLMs: $8k$.
@@ 4.2.1. Design
The evaluation involved testing each model on RustRepoTrans using controlled prompts, with output correctness assessed through specific test cases. The experimental design is structured as follows:
Evaluation Process. Each selected LLM was tasked with translating code into Rust, specifically focusing on 375 tasks. For each translation task, a corresponding set of test cases was employed to evaluate the correctness of the generated outputs. To ensure a fair comparison, the same prompt was used for each model, carefully designed based on established best practices in code translation tasks, as illustrated in Fig. 6. These best practices have been shown to enhance translation accuracy in similar studies (Macedo et al., 2024b; Khan et al., 2024; Pan et al., 2024; Yang et al., 2024). The prompt includes elements such as Instruction for translation and Translation’s required information(including source code, target function signature and target function dependencies)
@@
  • DSR@k: The DSR@k (Debugging Success Rate@k) metric, proposed in prior work (Yan et al., 2023), evaluates whether the generated code successfully executes and produces the expected results (i.e.,passing all test cases) within $k$ rounds of debugging, as expressed in Equation 2. $S(i,k)=1$ if the $i^{th}$ code sample succeeds within $k$ attempts; otherwise $S(i,k)=0$. In this study, DSR@k metrics were calculated with $k=1$, allowing for one debugging attempt.

### v4

# Repository-level Context Code Translation Benchmark Targeting Rust

@@ Abstract.
Recent advancements in large language models (LLMs) have demonstrated impressive capabilities in code translation, typically evaluated using benchmarks like CodeTransOcean. However, these benchmarks fail to capture real-world complexities by focusing primarily on simple function-level translations and overlooking repository-level context (e.g., dependencies). Moreover, LLMs’ effectiveness in translating to newer, low-resource languages like Rust remains largely underexplored.
To address this gap, we introduce RustRepoTrans, the first repository-level context code translation benchmark, comprising 375 tasks translating into Rust from C, Java, and Python. Using this benchmark, we evaluate four state-of-the-art LLMs, analyzing their errors to assess limitations in complex translation scenarios. Among them, Claude-3.5 performs best with 43.5% Pass@1, excelling in both basic functionality and additional translation abilities, such as noise robustness and syntactical difference identification. However, even Claude-3.5 experiences a 30.8% performance drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context compared to previous benchmarks without such context. We also find that LLMs struggle with language differences in complex tasks, and dependencies further increase translation difficulty.
copyright: acmlicensedjournalyear: 2025
@@ 1. Introduction
Code translation, or code migration, is the process of converting a software project from one programming language to another (Pan et al., 2023), often driven by the need to adapt to different runtime environments, improve performance, or enhance security (nee, 2018, 2020b, 2020c, 2020d, 2020a; Thönes, 2015). The rise of languages like Rust (rus, 2024) and Cangjie (can, 2025) has increased demand for translation. However, migrating legacy code remains challenging due to semantic mismatches, differences in standard libraries, and the need to maintain functional equivalence (Mateus et al., 2023; Kontogiannis et al., 2010).
Large language models (LLMs) have shown promise in translating widely used languages like Java and Python (Pan et al., 2024, 2023), but their effectiveness in translating to newer, lower-resource languages like Rust has not been thoroughly explored. Rust, an emerging language known for its safety guarantees, is increasingly in demand for translating projects from other languages to enhance memory safety and reliability (Security, 2024). However, its distinct features and limited training data pose challenges for automated translation, potentially limiting existing models’ performance (rus, 2024).
Existing code translation benchmarks, while covering Rust to some extent, have several limitations in capturing real-world scenarios. Most of them focus on function-level translation pairs (Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023), often translating self-contained functions with minimal dependencies (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a). Even file-level benchmarks only provide limited context for making isolated functions executable. Furthermore, many datasets are sourced from online coding platforms or artificially created examples, which differ greatly from real-world development practices (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a; Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023). In contrast, real-world translation occurs at the repository level, where developers incrementally translate code while preserving the project’s architecture. This involves managing complex dependencies, cross-file references, and inconsistencies between large-scale projects, aspects that current benchmarks fail to capture.
To bridge this gap, we introduce RustRepoTrans, the first benchmark to evaluate code translation with repository-level context, specifically targeting Rust. Unlike function-level benchmarks that assess translation in isolation, RustRepoTrans captures the complexities of real-world migration by incorporating repository-level context, including dependencies, cross-file interactions, and architectural constraints. These elements are crucial for ensuring that translated code integrates seamlessly into existing projects rather than functioning as isolated snippets. Our benchmark comprises 375 curated translation tasks derived from real-world projects spanning C, Java, and Python to Rust, providing a more faithful assessment of LLMs’ translation capabilities. While we focus on Rust due to its growing adoption in systems programming and the inherent challenges of translating into it, our methodology is generalizable to other programming languages, paving the way for broader repository-level evaluation.
@@
Building on RustRepoTrans, we assess state-of-the-art LLMs in translating code with complex repository-level context, analyzing challenges beyond existing benchmarks. We identify key factors affecting task difficulty and common errors, while also evaluating noise robustness, syntactical difference handling, and code simplicity during translation.
Our experiments show that LLMs struggle with repository-level code translation, with compilation errors reaching 94.8%, exposing the gap between current evaluations and real-world performance. Claude-3.5 performs best with 43.5% Pass@1, excelling in basic functionality and additional translation abilities like noise robustness and syntactical difference identification. However, even Claude-3.5 sees a 30.8% drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context. Dependency-related errors, including function and variable resolution issues, account for 61.9% of failures, highlighting the challenge of interconnected code. We also find that code complexity correlates with translation effectiveness. These insights reveal LLMs’ limitations in complex code migration, guiding future improvements in model design and evaluation.
@@
The contributions of this paper are as follows. Data and code are publicly available at  (rep, 2024).
@@
  • The first repository-level code translation benchmark, RustRepoTrans, with 375 tasks from C, Java, and Python to Rust, is introduced for realistic evaluation.
@@
  • Four state-of-the-art LLMs are evaluated on this benchmark, assessing their performance in real-world translation tasks.
  • LLM errors are categorized into 10 types, revealing limitations, especially in handling dependencies.
@@ 2.1. Code Translation
Code translation enhances software interoperability and facilitates legacy system updates, boosting productivity and reducing manual costs (Lu et al., 2021; Puri et al., 2021). Early methods relied on rule-based systems with handcrafted syntactic and statistical machine translation (SMT) models, improving accuracy by mapping frequent code patterns (Nguyen et al., 2013; Chen et al., 2018). Neural models, including RNNs, improved learning from data but struggled with complex syntax and idioms (Dong and Lapata, 2016; Yin and Neubig, 2017, 2018).
Despite advancements with LLMs, translation errors and adaptability issues persist (Zhu et al., 2022a; Rithy et al., 2022). Jiao et al. (Jiao et al., 2023) highlighted LLMs’ difficulty with complex translations, while Pan et al. (Pan et al., 2024) showed that context-rich prompts improve reliability by 5.5%. Yang et al. (Yang et al., 2024) and Macedo et al. (Macedo et al., 2024a) further refined prompt engineering for accuracy and consistency.
@@ 2.2. Code Translation Benchmarks
To address these gaps, some datasets, such as CoST (Zhu et al., 2022b) and CodeNet (Puri et al., 2021), have aimed to enhance language diversity and quality. However, issues like erroneous solutions and limited support for Rust persist. The latest benchmark, CodeTransOcean (Yan et al., 2023) and XCodeEval (Khan et al., 2024a), expands coverage to low-resource languages but still lacks adequate support and error analysis for Rust, hindering comprehensive research in this area.
@@
RustRepoTrans is the first repository-level context code translation benchmark focusing on Rust, designed to account for complex dependencies and derived from real-world projects, distinguishing it from previous benchmarks.
@@ 3. RustRepoTrans Benchmark
We introduce the benchmark format, construction method, and the resulting RustRepoTrans.
@@ 3.1. Benchmark Format
Each task in RustRepoTrans consists of a pair of functions with their dependencies and test cases, formatted as <source function, target function, target function dependencies, target function test cases> (Fig. 1). The function pairs represent functionally equivalent code snippets from source and target languages, with dependencies relevant to target function such as functions, data types, variables, and libraries. LLMs use the source function, target function signature, and associated dependencies to generate the target function, which is verified for correctness using the test cases.
@@ 3.2. Benchmark Construction Method
Migration Project Selection. In this step, we select projects that have been rewritten in Rust from other languages, specifically C, Java, and Python, due to their popularity and likelihood of having such rewritten versions. These projects are expected to contain functionally equivalent function pairs between their Rust and original language versions. We identify suitable projects by searching GitHub for terms like “Rust version” and “implemented in Rust”. focusing on larger projects (applying qualifier “size:>1000”) to ensure sufficient function pairs. After locating these projects, we verify their versions through documentation and code review. In this way, we identify pairs of source project and target project.
@@
Figure 2. Prompt for Identifying Equivalent Function Pairs
@@ 3.3. Resulting Benchmark
This process constructs RustRepoTrans with 375 repository-level translation tasks. Table 1 compares it with existing benchmarks. Similar in size to prior datasets, each task includes a manually verified ground truth translation and unit tests, achieving over 90% test coverage, ensuring high quality.
@@
Table 1. Comparison of Different Benchmarks for Code Translation
@@
RustRepoTrans Github Function Level 150.1 375 C, Java, Python
@@
Repository-level Dependency. RustRepoTrans focuses on code translation tasks with rich repository context, setting it apart from previous benchmarks. Unlike artificially constructed data or data from Q&A websites, real project data exhibits more complex dependency relationships, including function, data type, and variable dependencies, which are absent in existing datasets as shwon in Table 1, revealing that only RustRepoTrans includes these crucial elements. This inclusion makes RustRepoTrans a more realistic benchmark, suitable for evaluating LLMs that must account for intricate file-level interactions and contextual dependencies.
@@ 4.1. Experimental Setup
Model Selection. We selected four state-of-the-art LLMs that have been widely studied in recent code translation research or were recently released and have demonstrated strong performance on benchmark leaderboards (e.g.,SuperCLUE (Xu et al., 2023)). Table 2 provides an overview of the four LLMs used in our experiments, along with their release dates (Column “Time”) and model sizes (Column “Size”). As shown in Table 2, our study includes both open-source and closed-source LLMs, as well as models designed for general-purpose and code-specific applications. Due to resource constraints, we selected open-source models with sizes below 20B parameters. Note that our primary contribution is a new benchmark designed to provide a more challenging and realistic evaluation of LLMs’ code translation capabilities, closely simulating real-world development scenarios. Expanding the scope to include a larger and more diverse set of models is reserved for future work.
Implementation Details. For open-source models, we obtained and executed the released versions from their official repositories based on provided documentation with greedy sampling (Chen et al., 2015) as our generation strategy. All evaluations were conducted on an NVIDIA A800 80GB GPU. For closed-source LLMs, we accessed each model through its respective API interface (OpenAI, 2024; Claude, 2024) as of September 2024. To achieve results similar to greedy decoding, we set the “temperature” hyperparameter to 00.
@@ 4.2.1. Design
The evaluation involved testing each model on RustRepoTrans using controlled prompts, with output correctness assessed through specific test cases.
Evaluation Process. Each selected LLM was tasked with translating code into Rust, specifically focusing on 375 tasks. For each translation task, a corresponding set of test cases was employed to evaluate the correctness of the generated outputs. To ensure a fair comparison, the same prompt was used for each model, carefully designed based on established best practices in code translation tasks, as illustrated in Fig. 3. These best practices have been shown to enhance translation accuracy in similar studies (Macedo et al., 2024b; Khan et al., 2024b; Pan et al., 2024; Yang et al., 2024). The prompt includes elements such as Instruction for translation and Translation’s required information(including source code, target function signature and target function dependencies)
@@
  • DSR@k: The DSR@k (Debugging Success Rate@k) metric, proposed in previous work (Yan et al., 2023), evaluates whether the generated code successfully executes and produces the expected results (i.e.,passing all test cases) within $k$ rounds of debugging, as expressed in Equation 2. $S(i,k)=1$ if the $i^{th}$ code sample succeeds within $k$ attempts; otherwise $S(i,k)=0$. In this study, DSR@k metrics were calculated with $k=1$, allowing for one debugging attempt.